OpenStudy (anonymous):

n^2-16n+64 is prime right?

hartnn (hartnn):

it can be factored

OpenStudy (anonymous):

i tried using foil method but it didnt work

hartnn (hartnn):

tell me two numbers whose product is 64 and sum is -16

OpenStudy (anonymous):

\[\Delta=16^2-4\times 64=0\]yeah it is not prime

OpenStudy (anonymous):

so the way it is factored is (n-8)^2

hartnn (hartnn):

yup

OpenStudy (anonymous):

ok thanks alot